OpenStudy (anonymous):

secx= sqrt of 5 and0

OpenStudy (ranga):

\[\Large cos(x) = \frac{1}{sec(x)} = \frac{1}{\sqrt(5)} \]\[\Large \cos(2A) = 1 - 2\sin^{2}(A)\]Therefore, \[\Large cos(x) = 1 - 2\sin^{2}(\frac{x}{2})\]Substitute and solve for sin(x/2). Once you find sin(x/2) you can find cos(x/2) because sin^2(x/2) + cos^2(x/2) = 1. Then you can find tan(x/2) = sin(x/2) / cos(x/2).